OpenStudy (anonymous):

i need to answer this question to graduate! someone please help :'( Define the function N by the relations N(Alex) = 10, N(Ben) = 5, N(Cindy) = 0, N(Daisy) = 23, and N(Eric) = 10. N(Ben) + N(Eric) = _____.

OpenStudy (anonymous):

add

OpenStudy (anonymous):

add \(5+10\)

OpenStudy (anonymous):

but how do u put it in numeric form??

OpenStudy (anonymous):

N=15?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what equals 15?

OpenStudy (anonymous):

the answer is not N = anything the answer is N(Ben) + N(Eric) = 5+10=15

OpenStudy (anonymous):

that was the question right? question is N(Ben) + N(Eric) = _____.

OpenStudy (anonymous):

when i put that then it tells me that the format is wrong and i cant submit it :/

OpenStudy (anonymous):

it is asking you to add

OpenStudy (anonymous):

did you write 15?

OpenStudy (anonymous):

that worked! THANKS AHAH :)

OpenStudy (anonymous):

yw